OpenStudy (anonymous):

Which of the following is equivalent to |x − 12| = 6? A. x − 12 = 6 B. x − 12 = -6 C. x − 12 = 6 or x − 12 = -6 D. none of the above

sam (.sam.):

|x − 12| = 6 \[(x − 12) = \pm 6\]

sam (.sam.):

x-12=6 x=-6 x-12=-6 x=-18